OpenStudy (anonymous):

someone help pls, medal and fan! Without solving, how many solutions does the following equation have? Justify your answer in your written work. \begin{cases} y = -2x-17 \\ y = x^{2}+10x+22 \end{cases}

OpenStudy (anonymous):

there will be two solutions. If you let y=y -2x-17= x^2 +10x +22 then 0= x^2 + 12x +5 this is a quadratic equation. Quadratic equations generate 2 solutions.

OpenStudy (anonymous):

wait so that's the final answer?

OpenStudy (anonymous):

Answer: 2 solutions. What I wrote was the written work they asked for.

OpenStudy (anonymous):

thanksss! :D

OpenStudy (anonymous):

You're welcome :)